OpenStudy (anonymous):

The circle has a circumference of 22pi inches then what would its radius be * i need radius to find an arc length

jimthompson5910 (jim_thompson5910):

Circumference = 2*pi*radius C = 2pi*r C/(2pi) = r r = C/(2pi)

jimthompson5910 (jim_thompson5910):

Now plug in C = 22pi and simplify to find r

OpenStudy (anonymous):

well 22pi=69.08 would i sq that ?

jimthompson5910 (jim_thompson5910):

No need to approximate it, just plug it in and simplify r = C/(2pi) r = (22pi)/(2pi) r = 22/2 r = 11